OpenStudy (anonymous):

Compute: 1 3/4 + 4 3/5 =?

OpenStudy (anonymous):

\[1 \frac{ 3 }{ 4 } + 4 \frac{ 3 }{ 5 } \]

OpenStudy (anonymous):

Find a common denominator. 4 and 5 both multiply into what?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Good. I haven't done this in awhile but i'm pretty sure you have to have just a fraction, I'm not sure how to explain this really, so let me figure out the problem on my own then I'll try to see if I can explain. \[1\frac{ 3 }{ 4 } + 4\frac{ 3 }{ 5 }\] \[\frac{ 7 }{ 4 } + \frac{ 23 }{ 5 }\] Find a common denominator. Both 4 and 5 go into 20, so use 20 as the common denominator. When changing the denominator you have to multiply the top by the same number you did to the bottom. So you get: \[\frac{ 35 }{ 20 } + \frac{ 92 }{ 20 }\] Add. \[\frac{ 35 }{ 20 } + \frac{ 92 }{ 20 } = \frac{ 127 }{ 20 }\] Simplify. How many times does 20 go into 127? Six times with 7 left over. Write the equation. \[6\frac{ 7 }{ 20 }\] ^There's your answer.